OpenStudy (anonymous):

hat is the value of the variable that makes the statement true? b + (–2) + 8 = 0 A. –10 B. –6 C. 6 D. 10

OpenStudy (anonymous):

Your answer is B. -6 because -6+(-2)= -8+8=0.

OpenStudy (anonymous):

\[b+(-2)+8=0 <=> b+6=0 <=> b=?\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yep